# MAAdvisor: Zero-Shot Index Advisor using Multi-Agent LLMs

### v1

@@ 4. MAAdvisor Pipeline Overview
MAAdvisor
4.1. Workload Representation
Motivation: To enable zero-shot LLM reasoning effectively and efficiently, two main challenges lie within the information for index recommendation, namely long input context and lack of domain knowledge. First, the database information, like schema, workloads, and query plans, can be extremely long in real-world applications. For example, a workload consisting of 83 queries from the TPC-DS benchmark alone can take 36518 tokens’ length in GPT-4’s input. Not only do smaller LLMs may not have the ability to handle such long input sequences, but the cost and hallucination (Ji et al., 2023; Zhang et al., 2023) for larger LLMs dealing with long contexts can also be considerably large. Therefore, a new representation is required to summarize the database-related information into effective and precise LLM inputs. Second, although previous work (Wei et al., 2023) has proven LLMs’ strong generalization and reasoning ability given inputs similar or related to their training data, general-purpose LLMs always have insufficient training data related to data management and database indexing. Therefore, effective workload representation is vital for solving the index recommendation problem with LLMs.
Unfortunately, the workload representation techniques from existing methods are not fit for LLMs. The learning-based methods usually encode the database and workload information into embedded features (Lan et al., 2020; Wang et al., 2024b). However, such features are too abstract for LLMs to understand and use as inputs. The prompt-based method (Zhao et al., 2025) also attempts to address this problem by extracting index-related information in the workload for the LLMs. However, the information is yet stored in densely structured nested dictionaries. It is difficult for LLMs to locate the information needed, and hallucination may be triggered according to our experiments in Section 7.2. Therefore, a workload representation technique that is able to both shorten the LLM input length and preserve the critical information is required for LLMs to perform the index recommendation task.
Overview: To address the challenges above, we reformulate and summarize the inputs in a textual way that LLMs can easily understand and use. The information from workload, database schema, and optimiser that is relevant to index recommendation is extracted, merged, and grouped by the indexable column candidates. Given all the column candidates and their relevant information affecting index recommendation, the problem of using LLMs to construct indexes from raw workload can be simplified into selecting and combining columns from existing candidate choices. At the same time, by merging inputs on a per-column basis, the LLM’s prompt length is capped at the database’s total column count. In other words, regardless of how many queries appear in the workload, the set of candidate columns can never surpass the number of columns defined in the schema.
4.2. Multi-agent Pipeline
Motivation: To explore the capability of LLMs to solve the index recommendation task, we first conduct a pioneering experiment on the TPC-DS benchmark. We generate 100 small workloads, with each consisting of 3 to 10 random queries querying the TPC-DS database, and set a 2 GB storage budget. The results of directly letting GPT-4o (Achiam and et al., 2023) select indexes in a zero-shot manner given column candidates, and the heuristic-based baseline Extend (Schlosser et al., 2019) are presented in Table 2. It is observed that GPT-4o underperforms Extend by 25.42 seconds on average. Although introducing GPT-4o increases the scalability of the index recommendation, it demonstrates that the problem is yet too complicated for a single call of LLM to solve in zero-shot. Similar findings have been reported in (Zhao et al., 2025) that even with one-shot demonstration, LLMs still fail to outperform existing methods. However, when we remove the regressing indexes from the workloads and only keep the improving ones, it is observed that GPT-4o can outperform Extend. It demonstrates LLM’s potential in index recommendation and also justifies that an efficient and well-designed framework powered by LLM can successfully address the problem. Pioneer experiments also reveal that GPT-4o’s index recommendations fluctuate considerably, highlighting the need for a budget-aware planning pipeline to ensure the LLM produces consistent outputs while adhering to budget constraints.
Overview: To address the challenges above, we propose to simplify the index recommendation problem using a hierarchical multi-agent pipeline with budget-aware planning. At the global level, the multi-agent pipeline reflectively decomposes the problem into LLM-solvable sub-steps, including selecting, revising, or combining column candidates. At the local level, each sub-step is handled by a specialized local agent embedded with LLM. According to language model theory (Brown et al., 2020), LLMs produce more accurate and consistent outputs when handling simpler tasks. By breaking down the complex problem into smaller, well-defined sub-tasks and delegating them to specialized LLM agents, our multi-agent pipeline achieves both stability and efficiency in index recommendation. Moreover, by estimating the storage consumption at each sub-step, subsequent steps are planned accordingly, enabling budget-aware index recommendation.
@@ 5. Workload Representation
In this section, we will introduce our workload representation technique. In Section 5.1, we will first introduce the information we extract as a representation for each query in the workload. Then in Section 5.2, we will explain how we reformulate and merge the information in the whole workload into column candidates.
5.1. Query Information for Index Selection
To fully leverage the reasoning capabilities of LLMs, the information input should be concise and readily interpretable. However, the data extracted directly from SQL queries is often highly structured and domain-specific. Therefore, it is necessary to extract and reformulate the query-related information for the LLMs to interpret and process effectively. Building on insights from prior studies (Zhao et al., 2023, 2025; Siddiqui et al., 2022) and our preliminary experiments, we summarize key principles for extracting query-level information that are most relevant and effective for index recommendation with LLM:
Principle 1: To perform zero-shot inference, information that varies with databases and does not significantly affect index selections, for example, the data values in join conditions, predicates, and database knobs, can be excluded from the input;
Principle 2: Only columns that appear in any of the workload queries need to be considered as candidates to construct indexes, which will be called ‘indexable columns’.
Based on the principles above, the information we extract from a query representing each indexable column is as follows:
Table and column name: Table and column names are the key attributes for indexes. In addition, the semantics of the table and column names can usually offer the LLMs additional information during index recommendation. For simplification, the format ‘TableName.ColumnName’ will be used in the column candidate for LLMs to choose during the selection phase;
@@ 5.1. Query Information for Index Selection
Estimated column cardinality: As one of the key attributes affecting index selection, cardinality represents the uniqueness of data values contained in a column. Generally, indexes on higher cardinality columns allow easier access to specific rows. However, indexes on low cardinality columns may also help in specific situations, for example, the column value queried is extremely infrequent compared to the other values in the same column;
@@
Estimated column utility: Apart from the information above, an estimation of the potential improvement from a column index is vital for telling if a column candidate is worth selecting (Siddiqui et al., 2022). We define the utility of a potential index on a query as follows:
@@
(Index Utility): Given an index to create and a query , the utility of creating the index for is defined by query total cost without and with index per unit of storage cost:
where the database optimizer uses to estimate the total time cost of the query , and uses to estimate the storage cost of the index .
@@
5.2. Column Candidate Generation
In Section 5.1, we outlined the information required from each query for the index recommendation task. However, there still exist considerable duplications in information derived from each column. For a workload with indexable columns, where each column appears in queries on average, a total of pieces of information are collected. To enable the LLM to provide effective selections, this information must be consolidated so that each candidate item corresponds to a distinct column. It is observed that, among the information collected for a column, only the column operator type and the estimated column utility vary across queries. The column operator type can easily be merged by recording all the operators in a list. Consequently, an algorithm merging the estimated column utility for all queries in the workload can effectively eliminate duplicate column candidates.
Since the goal of index recommendation is to improve workload performance, the index that reduces the execution cost more should be considered as more beneficial. Therefore, we merge the estimated column utility for the same column in different queries by aggregating their utility values. If a column has a higher utility, it means that selecting the column as an index will result in an estimated higher workload performance gain. For the same workload above, there will be exactly column candidates left for the indexable columns after the merging. With the utility merged by the columns, each column candidate is represented as a dictionary of information so that the LLMs can easily access and read. It is also important to note that the merging also benefits the LLM efficiency, especially for larger workloads, by largely reducing the input length.
agents
@@ 6. Multi-agent Pipeline
In this section, we will introduce our multi-agent pipeline in which the index recommendation problem is decomposed into sub-steps for agents to solve. As shown in Figure 2, the pipeline consists of two types of agents: local agents and global agents. In Section 6.1, we will first introduce the three local agents that will select, combine, or revise the index recommendation in each sub-step. We will then explain our design for the two global agents planning each sub-step and reflecting on the plans in Section 6.2. In Section 6.3, we will illustrate the whole multi-agent pipeline on how we let the agents cooperate and communicate to perform index recommendation.
6.1. Local Agents
Selection agent: As the key function of our index advisor, the selection agent plays the role of selecting new columns from the column candidates given. The inputs for the selection agent include previous selected indexes, column candidates to select from, total storage budget, and the storage budget already used. In each single call of the selection agent, LLM is prompted with the inputs to select one new column into the current index recommendation as a single-column index.
Combination Agent: As noted earlier, the selection agent produces only single-column indexes. However, when multiple columns from the same table are frequently queried together, a composite index on those columns may deliver better performance than maintaining them as separate single-column indexes. To address this, we design a combination agent that can merge such indexes. To identify optimal composite indexes, we follow the approach in (Siddiqui et al., 2022), collecting potential rules for combining single-column indexes. As illustrated in Figure 3(b), these rules specify possible column combinations and their ordering based on query operations, which can be readily extracted from the queries themselves. In addition, we incorporate external knowledge from diverse sources such as the Internet, expert practices, and textbooks to fully leverage the LLM’s reasoning capabilities, thereby infusing domain expertise and facilitating effective index combination. Examples of external knowledge can also be found in Figure 3(c). Given the input of current selected indexes with operations, index combination rules, and external knowledge, the combination agent will decide if some indexes can be combined into a composite index and output the new index recommendation.
@@ 6.1. Local Agents
Revision Agent: The revision agent will be adopted to eliminate the potential duplicates incurred and index regressions caused by the current index recommendation. The duplicates can be easily removed by simple coding techniques. However, as addressed by the pioneer experiments in Table 2, index regression is one of the unsolved problems for both traditional and LLM-based index advisors. One of the reasons for the problem is the database optimizer’s estimation error for cardinality, which is used for evaluating the index’s potential benefits. Apart from the optimizer estimations, LLMs can also obtain cardinality-related information from the name and distribution of the columns. For example, the cardinality for the column ‘country_name’ is highly likely to be smaller than the column ‘order_id’. Therefore, we prompt the LLM to identify and remove columns from the selected index when a significant discrepancy exists between the estimated cardinality and the LLM’s analysis. To enhance the reasoning, we incorporate expert knowledge drawn from diverse sources such as common practices and online discussions. Insights of inefficient indexes are then distilled into expert experience rules, which guide the detection of regressive indexes. An example of such rules is shown in Figure 3(d).
@@
Regression indicator: Prior research (Ding et al., 2019; Shi et al., 2022) has attempted to address the index regression prediction problem using deep learning techniques. They label query plan pairs before and after index creation as either ‘improvement’ or ‘regression’, and train models on these labeled examples to forecast whether a given index will degrade query performance. However, these methods have limitations in handling zero-shot index recommendations at the workload level. Specifically, they focus only on query-level regressions rather than considering the impact on the entire workload. Moreover, their models require retraining for each new dataset, resulting in limited robustness and generalizability.
To address the aforementioned limitations, we devise a workload-level regression indicator that is trained on features invariant to specific databases and workloads. The overall architecture is depicted in Figure 4. The workload’s query plans, both with and without the indexes, are extracted as inputs for the indicator. Each operator in a plan is represented as a node, forming a binary tree for the entire query. Following the approach in (Zhao et al., 2022), we encode each query tree using tree-biased attention, where a node attends only to those within its sub-tree. The root node then serves as the vector representation of the entire query. The workload representation is constructed by concatenating the root vectors of all queries. Finally, given the pair of workload vectors before and after index creation, we employ an MLP classifier to predict potential index regressions from -1 to 1 based on the difference between the two representations. Each index will be assigned a predicted regressive score that the LLM can use as a reference.
@@
regression_indicator
6.2. Global Agents
@@
Planning agent: The planning agent plays an important role in the multi-agent pipeline for managing problem sub-steps and determining which local agents to deploy at each step. Given the column candidates, current indexes, total storage budget, and budget already used, the planning agent first evaluates the stop criteria. If these criteria are not met, it outputs an action for the current sub-step: ’Selection’, ’Combination’, or ’Revision’, each invoking a corresponding local agent. When the planning agent determines that the index selection is optimized or that no additional columns can be chosen within the budget, it triggers the ’Stop’ action to conclude the optimization.
@@
Reflection agent: In practical use, it has been observed that planning agent’s strategies for local agents can be inefficient at times, as it may suffer from hallucinations or become trapped in local optima. For instance, the planning agent may repeatedly revise the current index recommendation, invoke the combination agent when no viable combinations exist, or produce invalid actions due to hallucinations. To enhance the planning process, a reflection agent is introduced to evaluate the current path, to identify and correct errors. This helps the planning agent escape local optima and supports more efficient sub-step planning. Given the sequence of previous sub-step actions and the utility of current indexes, the reflection agent will offer suggestions for the next step’s planning. Example of possible suggestions can also be found in Figure 3(a).
6.3. Multi-Agent Pipeline

### v2

@@ 4. Framework Overview
AMAZe
4.1. Multi-agent Based Pipeline
Motivation. Although LLMs demonstrate strong potential for zero-shot generalization in index recommendation, it is hard to optimize the LLM prompts for each database schema. To investigate this limitation, we conduct a preliminary experiment on the TPC-DS benchmark. We construct 100 small workloads, each containing 3–10 randomly selected queries over the TPC-DS schema, under a fixed 2 GB storage budget. We compare the performance of GPT-4o (Achiam and et al., 2023), applied directly in a zero-shot manner with workload and query plan inputs, against the heuristic baseline Extend (Schlosser et al., 2019). Table 2 shows that GPT-4o underperforms Extend by an average of 15.8 % (25.42 seconds). While GPT-4o can generate reasonable index recommendations without additional training, its performance remains limited and inconsistent across complex workloads. These findings confirm that index recommendation is far too complex to be solved effectively with a single zero-shot LLM call, highlighting the need for a well-designed LLM-powered framework that can fully harness LLMs’ zero-shot generalization capabilities beyond direct prompting alone.
Inspired by the recent advancements in multi-agent LLM pipelines that divide a complex problem into sub-problems handled by specialized agents (Liu et al., 2025; Li et al., 2025; Jha et al., 2025), we aim to decompose the index recommendation task into multiple sub-tasks like index selection and combination. This design enhances scalability, interpretability, and efficiency, particularly well-suited for complex workloads that demand coordination, specialization, and iterative refinement.
Pipeline Overview. We introduce a solution based on a hierarchical multi-agent based index recommendation pipeline, illustrated in Figure 1. At the high-level, the pipeline reflectively decomposes the complex index recommendation task into a sequence of LLM-solvable sub-steps, such as selecting promising column candidates, revising suboptimal or redundant choices, and composing indexes when beneficial. At the low-level, each sub-step is handled by a specialized agent powered by an LLM, ensuring that decisions remain accurate and contextually appropriate. This design directly addresses the challenge that LLMs struggle with the full index recommendation problem in a single zero-shot call: the multi-agent pipeline excels at solving simpler, well-defined tasks with greater accuracy and consistency (Brown et al., 2020). By orchestrating multiple agents, our pipeline leverages the generalization power of LLMs while avoiding their weaknesses on overly complex reasoning.
4.2. Key Component Designs
Design Rationale. To develop an effective multi-agent framework, we must solve three critical design problems: (1) how to efficiently represent workloads for LLM agents, (2) how to incorporate storage budget constraints into the agent pipeline, and (3) how to prevent index regression that undermines recommendation quality. These design challenges are elaborated below:
(1) Workload Representation. Jointly encoding database schemas, workloads, and query execution plans often exceeds the context lenght limitations of LLMs—the TPC-DS workload with 83 queries requires over 36,000 tokens, surpassing most LLM’s context capacity and incurring substantial costs. Beyond computational constraints, long contexts reduce agent scalability and decision accuracy. Moreover, existing workload representations are unsuitable for agent-based systems. Learning-based methods encode information into embedded features (Lan et al., 2020; Wang et al., 2024b) that agents cannot interpret, while prompt-based approaches (Zhao et al., 2025) encode each query into a nested dictionary and append them as a list input that agents struggle to navigate effectively. These verbose, code-format representations misalign with how LLM agents process and reason about information. Therefore, we need to design a specialized workload representation technique that enables efficient agent communication and reasoning.
(2) Budget Constraints. Existing agent frameworks lack mechanisms to enforce storage constraints across multiple agents. Our preliminary experiments reveal that LLM does not always know when to stop the recommendation, demonstrating that agents operating without budget coordination produce inconsistent and resource-inefficient results. Thus, we must design budget-aware coordination mechanisms that enable agents to collaborate while respecting storage limitations.
(3) Regression Prevention. The index regression problem—where recommended indexes degrade query performance-is amplified in multi-agent settings due to three factors: suboptimal decisions by individual agents, LLM hallucinations (Ji et al., 2023; Zhang et al., 2023), and compounding errors across agent interactions. As shown in Table 2, when regressing indexes are removed, GPT-4o surpasses traditional baselines, indicating that not only regression prevention is crucial, but also optimal index recommendation is achievable by LLMs. However, general-purpose LLMs lack database domain knowledge (Wei et al., 2023), leading agents to make uninformed decisions that cascade through the pipeline. Existing agent frameworks provide no mechanisms to detect, prevent, or correct such regression patterns. Consequently, we need to design agent coordination strategies that incorporate database-specific validation, decompose complex decisions into reliable subtasks, and implement feedback mechanisms to prevent regression propagation.
Key designs. To address the aforementioned challenges, we propose three key designs that enable effective multi-agent coordination for index recommendation.
(1) Efficient Workload Representation. We design a novel textual workload representation that transforms the complex index recommendation problem into an LLM-compatible format. As illustrated in Figure 1, our approach extracts only index-relevant information and reorganizes it at the column level, merging all workload details associated with each database column to construct a compact set of indexable candidates. This design provides two critical advantages. First, it ensures efficiency and scalability by bounding the input length to the number of database columns rather than the number of queries—making the representation size independent of workload complexity. Second, it reduces hallucination by reframing index recommendation as a structured selection task where agents choose and combine predefined indexable columns, aligning the problem with LLMs’ decision-making strengths rather than requiring free-form generation.
(2) Budget-Aware Agent Planning. Our pipeline incorporates budget-aware coordination mechanisms that enable high-level agents to make resource-conscious decisions throughout the recommendation process. The planning agent receives the current storage consumption at its stage and communicates constraints to subsequent low-level agents, ensuring that the final recommendations remain feasible under real-world storage limitations. This index budget management prevents inefficient computations while maintaining the flexibility for agents to explore diverse optimization strategies within acceptable bounds.
(3) Effectiveness-Aware Agent Coordination. To prevent index regression, we design specialized low-level agents targeting the two primary causes of performance degradation: LLM hallucination and database optimizer estimation errors. As shown in Figure 1, our coordination strategy employs two complementary agents: The combination agent addresses optimizer estimation bias by merging individually suboptimal indexes into composite indexes that better capture multi-column access patterns, effectively combining inefficient recommendations into more beneficial ones. The revision agent eliminates LLM-generated errors, including hallucinated indexes that do not exist in the schema, redundant duplicates, and indexes that empirically degrade performance. Importantly, the cooperation between these two agents not only addresses hallucination and misestimation but also simulates a hybrid of ADD and DROP index recommendation heuristics, where new composite indexes can be constructed while regressive ones are systematically removed. This coordinated workflow enables the system to leverage the strengths of both heuristics, leading to a more robust, stable, and effective index recommendation pipeline than relying on any single mechanism.
agents
@@ 5. Workload Representation
In this section, we will introduce in detail our efficient workload representation. In Section 5.1, we will first introduce the information we extract as a representation for each query in the workload. Then in Section 5.2, we will explain how we further enhance pipeline efficiency by reformulating and merging the information in the whole workload into column candidates.
5.1. Query Information Extraction
To fully leverage the reasoning capabilities of LLMs, the input information should be concise and readily interpretable. However, data extracted directly from SQL queries is often highly structured and domain-specific, making it challenging for LLMs to process effectively. Therefore, it is necessary to extract and reformulate query-related information into a format that LLMs can interpret and process efficiently. Since we need to build index candidates derived from related columns, we first extract column information from workload queries. Building on insights from prior studies (Zhao et al., 2023, 2025; Siddiqui et al., 2022b) and our preliminary experiments, we identify that the essential column-related information includes:
Indexable column names: Table and column names are the key attributes for constructing indexes. We define each column that appears in the workload as an ‘indexable column’. Specifically, the format ‘TableName.ColumnName’ will be used in the column candidate for LLMs to choose during the selection phase;
@@ 5.1. Query Information Extraction
Estimated column cardinality: As one of the key attributes affecting index selection, cardinality represents the uniqueness of data values contained in a column. Generally, indexes on higher cardinality columns allow easier access to specific rows. However, indexes on low cardinality columns may also be beneficial in specific situations, such as when the queried column value is extremely infrequent compared to others in the same column;
@@
Estimated column utility: Apart from the information above, an estimation of the potential improvement from a column index is vital for telling if a column candidate is worth selecting (Siddiqui et al., 2022b). We define the utility of a potential index on a query as follows:
@@
(Index Utility): Given an index and a query , the utility of creating the index for is defined by query total cost without and with index per unit of storage cost:
where database optimisor uses to estimate the total time cost of the query , and uses to estimate the storage cost of the index .
@@
5.2. Workload Information Summarization
Motivation. Given a workload with queries, we could extract column candidates by individually processing each query and extracting an average of indexable columns per query. However, this approach is inefficient because many columns appear repeatedly across different queries, creating numerous duplicates. To enable the LLM to make effective and efficient index selections, we must consolidate this information so that each candidate corresponds to a unique column rather than being repeated across multiple queries.
To achieve this goal, we aggregate information for columns with the same name, ensuring that our final candidate set contains only unique columns. For most column attributes, consolidation is straightforward. The estimated column cardinality, column distribution, and estimated column storage remain constant for a given column regardless of which query references it, so we retain these values directly. However, two attributes vary across queries and require special handling:
(1) column operator types: We merge these by collecting all operators observed for a column across different queries into a single comprehensive list.
(2) estimated column utility: This requires a more principled aggregation approach. Since the goal of index recommendation is to enhance overall workload performance, columns that provide greater reductions in query execution cost should be prioritized. Therefore, we aggregate utility values across all queries that reference a given column. A higher aggregated utility indicates that indexing this column will likely yield greater overall performance improvement for the entire workload.
@@ 6. Multi-agent Pipeline
This section presents our hierarchical multi-agent pipeline for index recommendation, decomposing the problem into structured sub-steps handled by specialized agents. Section 6.1 introduces three low-level agents that select, combine, and revise indexes within individual sub-steps. Section 6.2 details two high-level agents responsible for planning and resource allocation under budget constraints. Section 6.3 demonstrates how these agents interact to achieve budget-aware planning and coordination.
6.1. Low-level Agents
Selection agent: As the key function of index advisor, the selection agent plays the role of selecting new columns from the column candidates given. As shown in Figure 2, the inputs for the selection agent include previous selected indexes, column candidates to select from, total storage budget, and the storage budget already used. Once the selection agent is called, LLM is prompted to select one new column into the current index recommendation as a single-column index.
Combination Agent: As noted earlier, the selection agent produces only single-column indexes, which are often insufficient when queries involve multiple columns from the same table. In such cases, composite indexes can provide greater efficiency than maintaining several separate single-column indexes. To address this, we introduce the combination agent, which merges selected single-column indexes into composite indexes that better capture multi-column access patterns. In doing so, it also mitigates optimizer estimation bias by transforming individually suboptimal indexes into beneficial composite ones.
To guide this process, we first adopt the approach in (Siddiqui et al., 2022b), which outlines rules for combining indexes, including valid column groupings and orderings inferred from query structures, as illustrated in Figure 3(b). Beyond combination rules, we further enhance the agent with external knowledge from sources such as textbooks, expert practices, and the Internet, enabling it to incorporate domain expertise into index construction. Examples of such knowledge are shown in Figure 3(c). As shown in Figure 2, given the current indexes, query operations, combination rules, and external knowledge, the combination agent determines whether certain indexes should be merged, and outputs updated recommendations.
@@ 6.1. Low-level Agents
regression_indicator
Revision Agent: The revision agent will be adopted to eliminate the hallucinations and potentially inefficient indexes in the current index recommendation, which may lead to index regression. First, a hallucination check will be conducted in the revision agent to identify and correct the wrong or non-existing index selections. Second, to mitigate the impact of potentially inaccurate estimations made by the database optimizer, a problem that is well documented even in traditional optimization pipelines, the LLM is guided by a three-fold justification mechanism that draws upon column information, expert experience and regression indications, as shown in Figure 2. Apart from the optimizer estimations, LLMs can also obtain cardinality-related column information from the name and distribution of the candidate columns selected. For example, the cardinality for the column ‘country_name’ is highly likely to be smaller than the column ‘order_id’. Therefore, we prompt the LLM to identify and remove columns from the selected index when a significant discrepancy exists between the estimated cardinality and the LLM’s analysis. To enhance LLM’s ability to tell a regressing index, we inject expert knowledge drawn from diverse sources such as common practices and online discussions. Insights of inefficient indexes are then distilled into expert experience rules, which guide the detection of regressive indexes. An example of such rules is shown in Figure 3(d).
@@
Regression indicator: Since prior learned index regression indicators (Ding et al., 2019; Shi et al., 2022) are dataset-specific and only detect query-level regressions, we propose a workload-level regression indicator which is trained using information invariant to databases and workloads. The structure of the model is presented in Figure 4. The workload’s query plans, both with and without the indexes, are extracted as inputs for the indicator. Each operator in a plan is represented as a node, forming a binary tree for the entire query. Following the approach in (Zhao et al., 2022), we encode each query tree using tree-biased attention, where a node attends only to those within its sub-tree. The root node then serves as the vector representation of the entire query. The workload representation is constructed by concatenating the root vectors of all queries. Finally, given the pair of workload vectors before and after index creation, we employ an MLP classifier to predict potential index regressions based on the difference between the two representations. Each index will be assigned a regressive score predicted that the LLM can use as a reference.
@@
6.2. High-level Agents
@@
Planning agent: The planning agent plays an important role in the multi-agent pipeline for managing problem sub-steps and determining which Low-level agents to deploy at each step. As shown in Figure 2, given the column candidates, current indexes, total storage budget, and budget already used, the planning agent first evaluates the stop criteria. If these criteria are not met, it outputs an action for the current sub-step: ‘Selection’, ‘Combination’, or ‘Revision’, each invoking a corresponding low-level agent. When the planning agent determines that the index selection is optimized or that no additional columns can be chosen within the budget, it triggers the ‘Stop’ action to conclude the optimization.
@@
Reflection agent: In practical use, it has been observed that planning agent’s strategies for low-level agents can be inefficient at times, as it may suffer from hallucinations or become trapped in local optima. For instance, the planning agent may repeatedly revise the current index recommendation, invoke the combination agent when no viable combinations exist, or produce invalid actions due to hallucinations. To enhance the planning process, a reflection agent is introduced to evaluate the current path, with the goal of identifying and correcting errors. This helps the planning agent escape local optima and supports more efficient sub-step planning. Given the sequence of previous sub-step actions and the utility of current indexes, the reflection agent will offer suggestions for the next step’s planning. Example of possible suggestions can also be found in Figure 3(a).
6.3. Multi-agent Planning and Coordination